Determine the mass(g) of calcium hydroxide in 2.55 L of solution, if its pH is 11.64 .
1
Expert's answer
2020-04-06T11:46:13-0400

pOH=14-pH=-lg[OH^-]

[OH^-]=-10^(-(14-11.64))=4.37*10^(-3) mol/L

Ca(OH)2=Ca^(2+)+2OH^-

[Ca^(2+)]=4.37*10^(-3)/2=2.18*10^(-3) mol/L

mol Ca^(2+)=2.55*2.18*10^(-3)=5.57*10^(-3)

m=5.57*10^(-3)mol*40.078g/mol=0.223g